OpenStudy (anonymous):

If W is the center of the circle below, and arc XY measures 52, what is the measure of angle xyz , in degrees?

OpenStudy (anonymous):



OpenStudy (nurali):

angle XWZ = 52° Assuming YZ is a diameter (since it passes thru the center W): Angle XWY = 180° - 52° = 128° [because of the adjacent angles rule] Now WX = WY = radius of circle So ΔWXY is isosceles So angle WXY = angle WYX = (1/2) [180° - 128°] = 26° and angle XYW = angle XYZ [actually the same angle] so angle XYZ = 26°

OpenStudy (imstuck):

Don't you mean what is the measure of angle XZY?

OpenStudy (anonymous):

 Angle XZY has the arc length 52 so knowing the relationship of an inscribed angle with its vertex on the circles outline you can find the angle easily \[\theta = \frac{ 1 }{ 2} arc\]